A drop of oil with a volume of 0.003 cubic mm has spread out over the surface of water in a
monomolecular layer and has covered 300 quadratic cm. Find the diameter of the molecules.
1
Expert's answer
2011-07-29T14:24:49-0400
V=0.003 mm2=3×10-15 m2
S=300 cm2=3×10-2 m2
It’s obvious that if we multiply the area covered with oil one molecule thick on the diameter of one molecule then we get the volume of oil :
Sd=V
Thus
d = V/S = (3×10-15 m2) / (3×10-2 m2 ) = 10-13 m = 0.0001 nm